OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5.

OpenStudy (radar):

Express this variation as a formula: y=k/x Where k is the constant of proportionality.

OpenStudy (radar):

Next find the value of k using the known situation -6=k/-2 We can see that k=12.

OpenStudy (radar):

We now solve the new condition: y=12/5=2.40

OpenStudy (radar):

Were you following the progression with understanding?

OpenStudy (anonymous):

Ahh I see now :) thank you.